Multiple Choice Question 67

Sunland Company’s high and low level of activity last year was 51000 units of product produced in May and 12000 units produced in November. Machine maintenance costs were $161800 in May and $52600 in November. Using the high-low method, determine an estimate of total maintenance cost for a month in which production is expected to be 36000 units.

$142800
$115050
$119800
$136050
Answer #1

$119800

Step-1:Calculation of variable cost per unit
Variable cost per unit = (a-b)/(x-y) Where,
= (161800-52600)/(51000-12000) a = Maintenance cost at highest level = $       1,61,800
= $       2.80 b = Maintenance cost at lowest level = $           52,600
x = High level of activity =               51,000
y = Low level of activity =               12,000
Step-2:Calculation of fixed cost
Level of activity a               51,000           12,000
Total costs b $       1,61,800 $       52,600
Variable cost per unit c $               2.80 $            2.80
Total Variable cost d=a*c $       1,42,800 $       33,600
Fixed Cost e=b-d $           19,000 $       19,000
Note:
Fixed cost is same at all level of activity.
Step-3:Calculation of total maintenance cost at 36000 units
Level of activity a 36000
Variable cost per unit b $               2.80
Total variable cost c=a*b $       1,00,800
Total fixed Cost d $           19,000
Total maintenance cost e=c+d $       1,19,800